2x - 3y = -1 | y = x - 1
1
Expert's answer
2013-01-15T09:47:28-0500
2x - 3y = -1
y = x& - 1

Substitute y=x-1 into the first equation:
2x-3(x-1)=-1
2x-3x+3=-1
-x=-1-3
x=4
Now we can find y=x-1=4-1=3